# DyCE: Dynamic Configurable Exiting

### v1

@@ VI-E Ensemble vs Individual Exits
Figure 6: Performance comparison between searched configurations and standalone exits on ResNet-50. MLP-AxB refers to An MLP of A layers with B neurons in all hidden layers.
VII Extensive Design and Usage
This paper demonstrates DyCE with the example of image classification. However, DyCE can be built in alternate ways and deployed to for more applications and more tasks.
VII-A Besides Image Classification
This paper uses the image classification task as an example. However, there should be no barriers to adapting DyCE to other tasks and hierarchical models. DyCE can be applied to any models that are dividable along the depth if we can design checkpoints to generate a candidate prediction in a compatible format as the original output and produce prediction confidence from the output. For example, Transformers[24] for natural language processing is stacked by dividable blocks, and we can also interpret the confidence from the largest value in its output vector. Therefore it should be possible to apply DyCE for Transformer networks.
VII-B Hierarchical multi-tasking
@@
VIII Conclusion
@@
Our experiments validate the effectiveness of the proposed method for image classification tasks employing CNNs. However, the concept of DyCE is not restricted to these domains and could potentially be extended to other models and applications.

### v3

@@ 6.5 Ensemble vs Individual Exits
Figure 7: Performance comparison between searched configurations and standalone exits on ResNet50. MLP-AxB refers to An MLP of A layers with B neurons in all hidden layers.
7 Limitations and Future Work
This paper demonstrates DyCE with the example of image classification. However, DyCE can be applied to more applications. In this section, we discuss its limitation and the potential of using DyCE in more cases.
7.1 Constraints and Limitations
DyCE is a highly flexible framework with notable advantages, but it also has some limitations. During deployment, DyCE requires a dynamic computational graph for inference, which is not well-supported by certain software libraries or hardware accelerators. This requirement may complicate deployment in some scenarios. Additionally, DyCE introduces extra parameters for exit candidates, and the size and number of exits might be restricted in practical implementations. These constraints are common to most dynamic models, and the community is actively seeking solutions. From a development perspective, DyCE assumes that intermediate features from pre-trained backbones are informative enough for prediction. This assumption simplifies system development since the backbone network can then remain unmodified. However, this approach may not achieve optimal performance, as the backbones were not originally designed for this purpose. Co-designing the exits alongside the backbone network could potentially yield better results.
7.2 Besides Image Classification
This paper demonstrates DyCE using an image classification task as an example. However, DyCE can theoretically be adapted to other tasks and models without major barriers. DyCE can be applied to any model that can be divided along its depth, provided that exits are designed to generate candidate predictions in a format compatible with the original output and produce confidence scores for those predictions.
To adapt DyCE to other data types, such as video, audio, or text, the backbone encoder should be replaced accordingly. Encoder architectures for these data types, like RNNs [31] and Transformers [32], are also composed of stacked, dividable blocks, making them suitable for DyCE. When adapting DyCE for tasks beyond classification, the exit component should be designed and trained specifically for the new task. For instance, the exit component for object detection should function as a detector, while for image-to-text tasks, it should act as a caption generator.
7.3 Hierarchical Multi-tasking
@@
7.4 Distributed Inference
The early-exit based model inference is done in a hierarchical manner. Since each exit can generate independent outputs, it is possible to partition the model on different devices for inference. DyCE can be used with model partitioning methods[33, 34] to deploy models on multiple devices. In this situation, DyCE can replace MACs with measured latency while generating configurations. The inference can exit with a very short latency if the initial early-exit network is enough to give an acceptable result. For complex events, the edge device can transmit that sample to fog or cloud nodes for more comprehensive inference. This scheme deals with most samples locally to provide real-time feedback but also tackles complex events by using large networks on remote servers. With DyCE, the overall performance of this heterogeneous system will be estimable and controllable.
8 Conclusion
@@
Our experiments validate the effectiveness of the proposed method for image classification tasks employing vision models. However, the concept of DyCE is not restricted to these domains and could potentially be extended to other models and applications.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT in order to improve the readablity. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.